OpenStudy (anonymous):

sin(π/4)=cos @dumbcow

OpenStudy (dumbcow):

sin(pi/4) = sqrt2/2 cos x = sqrt2/2 ---> x = pi/4

OpenStudy (anonymous):

sin(π/6)=cos

OpenStudy (dumbcow):

let me explain it this way  sin(pi/6) = y and cos A = y thus sin(pi/6) = cosA pi/6 +A = pi/2 --> A = pi/3

OpenStudy (anonymous):

I get it!

OpenStudy (anonymous):

thank you for sooo much help @dumbcow

OpenStudy (dumbcow):

your welcome

OpenStudy (anonymous):

so I'm always going to set it up like this? sin(pi/#) = cosA @dumbcow

OpenStudy (dumbcow):

yes as long as its in the form sin = cos in which case the 2 angles will always add up to pi/2

OpenStudy (anonymous):

what if its cos=sin @dumbcow ?

OpenStudy (dumbcow):

same thing, it doesnt matter if you flip everything x=5 is same as 5=x

OpenStudy (dumbcow):

cos(pi/#) = sin A where A is still the other angle in triangle pi/# + A = pi/2

OpenStudy (anonymous):

so its the same way but the cos and sin are in different spots?

OpenStudy (anonymous):

@dumbcow

OpenStudy (dumbcow):

yeah

OpenStudy (dumbcow):

fyi i need to log off now if have more questions, just tag me and i can look at it later

OpenStudy (anonymous):

ALRIGHT THANK YOU SOO MUCH!